# Connecting dust and outflows in AGN: the intriguing case of NGC 6860

### v2

@@ 2 Observational data and reduction
The Chandra HETG data were reduced using the Chandra Interactive Analysis of Observations (CIAO, Fruscione et al. 2006) v4.15 software and Calibration Database (CALDB) v4.10.7. The chandra_repro script of CIAO and its associated tools were used for the reduction of the data and production of the final grating products (PHA2 spectra, RMF and ARF response matrices).
@@ 3 Spectral fitting
For our modelling we use the software SPEctral X-ray and UV modelling and analysis, SPEX, version 3.06.01111http://doi.org/10.5281/zenodo.2419563 (Kaastra et al. 1996, 2018). To evaluate the goodness of our fit we adopt $C$-statistics which is denoted as $C_{\rm stat}$ in the rest of the manuscript. SPEX uses $C_{\rm stat}$ as described in Cash (1979), but with an addition which allows an estimate of the goodness of the fit (for details see Kaastra 2017). Also, in our analysis we use proto-Solar abundance units from Lodders & Palme (2009).
@@ 3.2 X-ray continuum modelling
The XMM-Newton and Chandra data were obtained at different epochs and therefore the continuum shape might be different due to variability intrinsic to the source. To take this variability into account, we use the sectors option in SPEX. With this option, each dataset is allocated to a specific sector. In this way, the continuum parameters (which are variable over time) for each dataset are treated independently and the normalisation of the pow component is allowed to vary freely, which gives a more accurate description of the continuum. The free parameters of our continuum model, listed in Table 2, are the normalisation of the pow and comt components, the photon index ($\Gamma$) of the pow and the scale of reflection in refl. The SED is plotted in Figure 1.
@@ 3.3 Photoionisation and spectral modelling of the ionised outflows
To adequately account for the absorption lines originating from various ionic species we need four ${\tt pion}$ components, with different ionisation parameters, $\xi$. The ionisation parameter is defined as $\rm\xi=L/(n_{H}\cdot r^{2})$, where where L is the luminosity of the ionising source over the 1–1000 Ryd band (13.6 eV to 13.6 keV) in $\rm erg\cdot s^{-1}$, $N_{\mathrm{H}}$ the hydrogen density in $\rm cm^{-3}$, and $r$ the distance between the photoionised gas and the ionising source in cm (Tarter et al. 1969, Krolik et al. 1981). For the different pion components we have as free parameters: the column density of the absorber, the ionisation parameters as well as the velocity shift and velocity broadening. The pion components are added until all the observed absorption features are effectively modeled, and no further statistical enhancement in the fit is achievable.
The highest ionisation component (Comp. A) of our fit, $\rm log\xi=3.4$, produces the absorption from highly ionised species such as S XIV, S XV, Fe XIX, Fe XX, Fe XXI, and improves our fit by $\Delta C_{stat}$=870. A second ${\tt pion}$ component (Comp. B) with lower ionisation parameter of $\rm log\xi=2.9$ fits the absorption lines of O VII, O VIII, Ne IX, Ne X and improves our fit by $\Delta C_{stat}$=530. A third ${\tt pion}$  component (Comp. C), $\rm log\xi=2.3$, is needed to improve the fit by $\Delta C_{stat}$=200. Finally we added a fourth weakly-ionized/neutral ${\tt pion}$  component (Comp. D) in the AGN. This component improves our fit by $\Delta C_{stat}$=1200 and contributes to the continuum absorption. Due to the lack of line detection, the velocity broadening ($\rm\sigma$) and velocity shift ($\rm z_{v}$) of this component are unconstrained. Therefore, we have fixed them at the default values, as they do not impact the fit. The best fit values of the pion components are presented in Table 3, and Figure 2 shows the best fit model on the data.
@@ 4.3 The nature and origin of dust
We can then calculate how far the outflowing gas would travel to reach the sputtering timescale: ${r\sim t_{\rm sput}\cdot z_{\rm v}}$, where $z_{\rm v}$ is the outflow velocity of pion #3 (Table 3). This assumes a simple steady flow. We thus find a distance of 83.6 pc. Therefore, if pion 3 is located at $\rm r<83.6$ pc, it would survive sputtering. Since our derived distance for pion #3 is $r<237$ pc (see Sect. 4.1), this means it is still feasible for dust to survive sputtering in this component.
@@ 5 Conclusions
We would like to thank Erin Kara for insightful discussions. This work is supported by NASA grant 80NSSC23K0490 for our XMM-Newton observation. This research has made use of data obtained from the Chandra Data Archive, and software provided by the Chandra X-ray Center (CXC). SRON is supported financially by NWO, the Netherlands Organization for Scientific Research. Support for this work was also provided by NASA through the Smithsonian Astrophysical Observatory (SAO) contract SV3-73016 to MIT for Support of the Chandra X-Ray Center (CXC) and Science Instruments. CXC is operated by SAO for and on behalf of NASA under contract NAS8-03060.

### v3

@@ 2 Observational data and reduction
The Chandra HETG data were reduced using the Chandra Interactive Analysis of Observations (CIAO, Fruscione et al. 2006) v4.15 software and Calibration Database (CALDB) v4.10.7. The chandra_repro script of CIAO and its associated tools were used for the reduction of the data and production of the final grating products (PHA2 spectra, RMF and ARF response matrices). To improve the signal-to-noise ratio and enhance our ability to detect absorption lines, we need to combine the Chandra observations. The data are divided into numerous short exposures, but after analyzing the overall flux variations between them, we find no significant evidence of variability. Additionally, the spectral shape of the continuum, as determined from our HETG fit, aligns well with the lower-energy shape observed with RGS. This suggests that NGC 6860 does not experience substantial variability during our observation period that would impact our modeling.
@@ 3 Spectral fitting
[…]
@@ 3.2 X-ray continuum modelling
The XMM-Newton and Chandra data were obtained at different epochs and therefore the continuum shape might be different due to variability intrinsic to the source. To take this variability into account, we use the sectors option in SPEX. With this option, each dataset is allocated to a specific sector. In this way, the continuum parameters (which are variable over time) for each dataset are treated independently and the normalisation of the pow component is allowed to vary freely, which gives a more accurate description of the continuum. The free parameters of our continuum model, listed in Table 2, are the normalisation of the pow and comt components, the photon index ($\Gamma$) of the pow and the scale of reflection in refl. The SED is plotted in Figure 1. The SED model we are using aligns well with the EPIC-pn data. We note that two etau components in SPEX are used to create the low-energy and high-energy exponential cut-off of the power-law component at reasonable energies. A high-energy cut-off has been applied at 300 keV and a low-energy cut-off at the Lyman edge (13.6 eV).
@@ 3.3 Photoionisation and spectral modelling of the ionised outflows
To adequately account for the absorption lines originating from various ionic species we need four ${\tt pion}$ components, with different ionisation parameters, $\xi$. The ionisation parameter is defined as $\rm\xi=L/n_{\mathrm{H}}\cdot r^{2}$, where L is the luminosity of the ionising source over the 1–1000 Ryd band (13.6 eV to 13.6 keV) in $\rm erg\cdot s^{-1}$, $n_{\mathrm{H}}$ the hydrogen density in $\rm cm^{-3}$, and $r$ the distance between the photoionised gas and the ionising source in cm (Tarter et al. 1969, Krolik et al. 1981). For the different pion components we have as free parameters: the column density of the absorber, the ionisation parameters as well as the velocity shift and velocity broadening. The remaining pion parameters have been kept at their default values, as specified in the user manual. [^2] https://spex-xray.github.io/spex-help/models/pion.html. In particular the density $n_{\mathrm{H}}$ of pion is left at its default value as our modeling is not dependent on density. The pion components are added until all the observed absorption features are effectively modeled, and no further statistical enhancement in the fit is achievable.
The highest ionisation component (Comp. A) of our fit, $\rm log\xi=3.4$, produces the absorption from highly ionised species such as S XIV, S XV, Fe XIX, Fe XX, Fe XXI, and improves our fit by $\Delta C_{stat}$=870. A second ${\tt pion}$ component (Comp. B) with lower ionisation parameter of $\rm log\xi=2.9$ fits the absorption lines of O VII, O VIII, Ne IX, Ne X and improves our fit by $\Delta C_{stat}$=530. A third ${\tt pion}$  component (Comp. C), $\rm log\xi=2.3$, is needed to improve the fit by $\Delta C_{stat}$=200. Finally we added a fourth weakly-ionized/neutral ${\tt pion}$  component (Comp. D) in the AGN. This component improves our fit by $\Delta C_{stat}$=1200 and contributes to the continuum absorption. Due to the lack of line detection, the velocity broadening ($\rm\sigma$) and velocity shift ($\rm v_{out}$) of this component are unconstrained. Therefore, we have fixed them at the default values, as they do not impact the fit. In our modeling the continuum first irradiates the highest-ionization pion component and reaches the lowest-ionization component last. The best fit values of the pion components are presented in Table 3, and Figure 2 shows the best fit model on the data. Finally, we examined potential variability in the pion components across different epochs using XMM-Newton and Chandra observations. However, we found no conclusive evidence of variability. We also examined our spectra for the presence of emission lines. However, these features appear too weak for a meaningful analysis. Typically, the strongest emission lines would be O VIII $Ly\alpha$ and the O VII forbidden line, but even these are barely detectable in our case. Our tests indicate that incorporating them into our modeling does not improve the fit, making it impractical to constrain their parameters.
@@ 4.3 The nature and origin of dust
We can then calculate how far the outflowing gas would travel to reach the sputtering timescale: ${r\sim t_{\rm sput}\cdot\rm v_{out}}$, where $\rm v_{out}$ is the outflow velocity of pion #3 (Table 3). This assumes a simple steady flow. We thus find a distance of 83.6 pc. Therefore, if pion 3 is located at $\rm r<83.6$ pc, it would survive sputtering. Since our derived distance for pion #3 is $r<237$ pc (see Sect. 4.1), this means it is still feasible for dust to survive sputtering in this component.
Laha et al. (2016) explore the dynamics of warm absorbers and their impact on the host galaxy’s environment. Assuming the observed outflow velocity of the warm absorber corresponds to the radial escape velocity ($\rm v_{out}$) at its distance from the central engine, the minimum launch radius ($\rm r_{min}$) can be estimated using the virial relationship:
$r_{min}=\frac{GM}{\rm v_{out}^{2}}$ (4)
where the black hole mass $M$ for NGC 6860 is ${4.0\times 10^{7}}$ $M_{\odot}$ (González-Martín & Vaughan 2012). Applying this formula, we find $\rm r_{min}=4\ pc$. Under this assumption, the pion component 3 is located within the range $2374<r<$ 237, where dust grains could potentially survive sputtering.
@@ 4.4 Comparison with other AGN with dusty outflows
The recent literature suggests that a major portion of the mid-infrared (MIR) emission in AGN originates from an extended polar component, indicating the presence of polar dust (Asmus 2019, and references therein). This polar dust is located at similar scales to the narrow-line region (NLR) of the AGN and likely forms a hollow, cone-like structure surrounding the NLR. Asmus (2019) studied nine nearby obscured Seyfert-type AGN using deep, subarcsecond-resolution MIR imaging and found that polar dust may be a ubiquitous feature of AGN structure. Therefore, if polar dust resides on scales comparable to the NLR, our results may imply that the observed dust in NGC 6860 is outflowing polar dust.
Finally, our study allows us to place limits on the distance of the absorbers and investigate the presence and location of dust grains in this AGN. Previously, NGC 6860 was studied by Winter & Mushotzky (2010), who analyzed Suzaku and XMM-Newton observations and identified a two-component warm ionized absorber in the soft X-ray spectrum. In our analysis, we identify three warm photoionised absorbers. This is likely due to differences in the coverage of high-resolution spectra (HETGS + RGS in this work), as well differences in modeling, which also impact the derived parameters.
4.5 The physical origin of the ionised outflows
One of the longstanding enigmas in AGN research is the nature and origin of ionised outflows (warm absorbers). Some studies have linked them to the presence of dust grains in AGN. Mizumoto et al. (2019) suggested that dust-driven winds may play a significant role in accelerating warm absorbers. Later on, Ogawa et al. (2022) investigated the ionization structure of the radiation-driven fountain model in a low-mass AGN (Wada et al. 2016). In this model, outflows are primarily driven by radiation pressure on dust and the thermal energy of gas resulting from X-ray heating, naturally forming a geometrically thick, torus-like structure. Their results suggest that additional components, originating closer to the SMBH than the torus—such as line-driven winds from the accretion disk—are likely crucial contributors to the formation of warm absorbers in AGN. Our results on NGC 6860 support this scenario, as we find that the dust in this AGN likely originates from the NLR and is present in the warm absorber.
The importance of radiation pressure on dusty gas as a source of AGN feedback has also been discussed in the literature (e.g. Ishibashi & Fabian 2016; He et al. 2022). Recently, Ishibashi et al. (2024) studied broad absorption line (BAL) outflows which are commonly detected in AGN but their driving mechanism is poorly understood. They concluded that dust-driven BAL outflows may provide a significant contribution to AGN feedback on galactic scales. In NGC 6860, we observe that the dust likely originates from the NLR and possibly the host galaxy, suggesting that the ionized dusty outflows could serve as a mechanism driving feedback within the host galaxy.
@@ 5 Conclusions
We would like to thank the referee for the suggestions that helped to improve the clarity of this paper. IP would like to thank Erin Kara for insightful discussions, and Jelle de Plaa for help with SPEX. This work is supported by NASA grant 80NSSC23K0490 for our XMM-Newton observation. This research has made use of data obtained from the Chandra Data Archive, and software provided by the Chandra X-ray Center (CXC). SRON is supported financially by NWO, the Netherlands Organization for Scientific Research. Support for this work was also provided by NASA through the Smithsonian Astrophysical Observatory (SAO) contract SV3-73016 to MIT for Support of the Chandra X-Ray Center (CXC) and Science Instruments. CXC is operated by SAO for and on behalf of NASA under contract NAS8-03060.
This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in  https://doi.org/10.25574/cdc.327 (catalog DOI: https://doi.org/10.25574/cdc.327).